Predetermined Factory Overhead Rate
Poehling Medical Center has a single operating room that is used by local physicians to perform surgical procedures. The cost of using the operating room is accumulated by each patient procedure and includes the direct materials costs (drugs and medical devices), physician surgical time, and operating room overhead. On January 1 of the current year, the annual operating room overhead is estimated to be:
Disposable supplies $299,600
Depreciation expense 75,000
Utilities 32,000
Nurse salaries 278,500
Technician wages 126,900
Total operating room overhead $812,000
The overhead costs will be assigned to procedures based on the number of surgical room hours. Poehling Medical Center expects to use the operating room an average of eight hours per day, seven days per week. In addition, the operating room will be shut down two weeks per year for general repairs.
a. Calculate the estimated number of operating room hours for the year.
hours
b. Determine the predetermined operating room overhead rate for the year.
$ per hour
c. Bill Harris had a five-hour procedure on January 22. How much operating room overhead would be charged to his procedure, using the rate determined in part (b)?
$
d. During January, the operating room was used 240 hours. The actual overhead costs incurred for January were $67,250. Determine the overapplied operating overhead or underapplied operating overhead for the period.
$
Answer #1

Estimated annual overhead = $ 812,000

a. Estimated total surgical hours = 2800 hours (8 x 7 x 50 )

b. Predetermined operating room overhead per hour = 812,000/2800 = 290 per surgical hr

c. Operating overhead = 5 x $290 = $1450

d. Actual overhead =$ 67,250

Applied overhead = 240*290 = $ 69,600

Overapplied operating room overhead = 69,600-67,250 = $ 2350